OpenStudy (anonymous):

List all the subsets of {1}.

OpenStudy (zehanz):

Remember that every set is also a subset of itself. Also the empty set is a subset of every set... There is not much left to think about now!

OpenStudy (anonymous):

so it is 1

OpenStudy (zehanz):

Hmm, (...counting...) no, it is 2!

OpenStudy (anonymous):

oh so it is 1 and 2 or just 2

OpenStudy (zehanz):

Listing of all the subsets of {1}: {1} The emptyset I think this listing is the answer.

OpenStudy (anonymous):

so {{}, {1}}

OpenStudy (zehanz):

Oh, I understand now, you are not supposed to just mention all the subsets, you have to give the set of all subsets of {1}. This is indeed:\[\left\{ \emptyset, \left\{ 1 \right\} \right\}\]

OpenStudy (zehanz):

Did you now that the number of subsets of a set is 2^(no of elements in the set)?

OpenStudy (anonymous):

No. my teacher put this on a study guide for a test and didnt teach it

OpenStudy (zehanz):

Impress/embarass him/her by asking this as a question ;) Another thing: the no of elements in a set is called the Cardinal number. Cardinal number of A={1, 2, 3} is 3. There are 8 subsets of A, because 2^3 = 8. The set of all subsets of A is:\[\left\{ \emptyset, \left\{ 1 \right\} , \left\{ 2 \right\},\left\{ 3 \right\},\left\{ 1,2 \right\},\left\{ 1,3 \right\},\left\{ 2,3 \right\},\left\{ 1,2,3 \right\}\right\}\]